For a given gas the coefficient of viscosity is 1.9 × 10⁻⁵ N s m⁻²

and the diffusion coefficient is 1.2 × 10⁻⁵ m² s⁻¹. Calculate the density and mean free path at average molecular velocity of 380 m s⁻¹.


1
Expert's answer
2021-04-13T07:02:50-0400

Answer

Density is given by

"\rho =\frac{\eta}{D}\\=\frac{1.9\times10^{-5}}{1.2 \times10^{-5} }\\=1.58kg/m^3"

Mean free path

"\lambda=\frac{1}{v_a}\\=\frac{1}{380}\\=2.63\times10^{-3}m"